OpenStudy (anonymous):

Find the area of a circle whose circumference is 31.4 centimeters.

OpenStudy (anonymous):

Okay, first step - what is the formula for circumference?

OpenStudy (anonymous):

31.4?

OpenStudy (anonymous):

31.4 is the value of the circumference given to you. Do you know the formula used to calculate circumference?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

Is it gonna be 3.14

OpenStudy (nirmalnema):

2 *pi *r = 31.4 r=31.4/(2*3.14) r=5 area = pi r*r =3.14*5*5 =78.5

OpenStudy (anonymous):

Well, the formula for the circumference of a circle is $$ 2 \pi r $$